# Seek in the Dark: Reasoning via Test-Time Instance-Level Policy Gradient in Latent Space

### v1

@@ 3.2 State-of-the-art Test-time Reasoning Performance
The main experimental results are presented in Table 1 and Table 2.
@@
As demonstrated in Table 2, using LLaMA3.1-8B-Instruct as the backbone, our method outperforms all baseline approaches across both the GSM8K and MATH-500 datasets. Specifically, on the GSM8K dataset, our method achieves an improvement of 14.6% points over the CoT baseline, and a 7.0%-point gain on the MATH-500 dataset. In comparison to the training-free BoN approach, our method yields improvements of 7.7% points and 3.4% points on GSM8K and MATH-500, respectively. When compared to the SFT (Magpie 25K) Xu et al. (2025a), the minimum improvement across both datasets is 12.9% points. Moreover, our approach surpasses the state-of-the-art self-reward RL-based method, Genius Xu et al. (2025a), by 5.4% points on GSM8K and 20.0 points on MATH-500. Notably, it also outperforms iCoT Deng et al. (2024)—a latent method trained on an augmented GSM8K dataset—by 36.6% points on the GSM8K test set. Even when compared to reinforcement learning with verifiable feedback methods, like SimpleRL-Zoo Zeng et al. (2025), which was trained on the base version of the model, our method exceeds it by an average of 9.1% points across both datasets. These results underscore the effectiveness of our approach in tackling reasoning tasks.
@@
AIME2024 results are listed in Table 1. AIME2024 is a highly challenging mathematics benchmark consisting of only 30 problems, designed to evaluate advanced reasoning capabilities. Experimental results demonstrate the power and potential of our method in complex settings. As shown in Table 1, our approach consistently outperforms all baselines, achieving an average improvement of 4.73% points over CoT across all model families and prompt configurations. Notably, even weaker models such as Mistral and Qwen2.5-1.5B-Instruct benefit from our method, with average gains of 1.67% and 3.33% points compared with CoT, respectively. The largest improvement is observed on AIME2024 (prompt 1), where our method surpasses the CoT baseline by 13.33% points and outperforms BoN by 6.67% points using Qwen2.5-7B-Instruct.
@@
As presented in Section 3, the generalizability of our approach with respect to backbone models can be examined along two dimensions: (1) Model Families: LatentSeek demonstrates superior performance across multiple model families. Specifically, it achieves the best results when using backbones from the Qwen2, Qwen2.5, and LLaMA3.1 series. Additionally, it ranks first in four out of six evaluation settings and second in one when employing a relatively weaker backbone from the Mistral family. The suboptimal performance of the Mistral-based model is attributed to its limited capacity to function effectively as a reward model. (2) Model Size: At the 1.5B parameter scale, our method consistently outperforms all baseline models across diverse datasets and prompt types, with a particularly notable gain of 28.89% points over the CoT baseline on the GSM8K dataset (Prompt 2). At the 7–8B scale, models based on Qwen2, Qwen2.5, and LLaMA3.1 consistently surpass all baselines. In particular, Qwen2 yields an average improvement of 11.75% points over BoN on GSM8K. At the 14B scale, our approach continues to outperform all baselines, achieving a significant 6.67%-point improvement on the AIME2024 dataset. These results provide robust evidence of our method’s generalizability across diverse model families and scales.
@@ 3.3 Ideal Experiment: Perfect Sparse Reward Model
To further evaluate the effectiveness of exploration within the latent space, we conduct experiments using a Perfect Sparse Reward Model (PSRM). In this experimental setup, a reward value of 0 is assigned exclusively when the generated final answer exactly matches the ground truth. In all other cases, a reward of $-1$ is given. The results of these experiments are presented in Table 1. Despite the limited feedback provided by this sparse reward signal—particularly in contrast to the self-reward model—the PSRM demonstrates notable performance gains. Specifically, it achieves an average improvement of 19.12% score over the CoT method and surpasses the self-reward version by an average of 12.57% score. These results highlight the potential of LatentSeek, even when guided by sparse reward.
@@ 3.4 Test-Time Scaling: scaling up the iteration of LatentSeek
In addition to increasing the number of generated tokens at test time Muennighoff et al. (2025), Snell et al. (2025), we propose an alternative approach to test-time scaling: increasing the number of LatentSeek iterations. The relationship between model performance on GSM8K Cobbe et al. (2021) and the number of iterations is illustrated in Figure 2. As shown, the ideal reward model yields a consistently improving trend and outperforms the self-reward method across all model backbones, suggesting that test-time scaling can be achieved without necessitating a dense reward function in our setting. Conversely, the self-reward method exhibits rapid initial gains followed by a plateau, attributable to the limitations in the reward model’s accuracy. Notably, with an increased number of iterations, performance eventually surpasses this plateau and continues to improve. These findings indicate that, given an appropriate reward model, searching through the latent space offers a promising new direction for test-time scaling.
@@
Following the methodology proposed by Liu et al. (2025), we conduct extreme scaling experiments using the Qwen2.5-1.5B-Instruct model Team (2024a). In contrast to Liu et al. (2025), which leverages Process Reward Models (PRM), our approach is specifically tailored to align with Outcome Reward Models (ORM). Given the current limitations in ORM design (see Appendix F) and the inherent challenges of self-rewarding in small-scale models, we investigate the effectiveness of this novel test-time scaling strategy through the use of the PSRM. As shown in Table 3, Qwen2.5-1.5B-Instruct achieves a 14% point performance gain over GPT-4o on the AIME2024 benchmark. Furthermore, on the MATH-500 dataset, it attains the highest overall accuracy among all evaluated models, with a marginal 2.7% point gap behind o1-preview. Remarkably, the model requires only an average of 61.8 iterations on MATH-500 to reach the performance, outperforming a 256 compute consumption of explicit space scaling methods in terms of efficiency. This implies that, under an appropriate reward model, the latent space represents a more efficient option for test-time scaling compared to the explicit space.
@@ 3.5 Algorithmic Statistics
Table 4 reports two key statistical features: (1) the ratio of final answer tokens to the original CoT tokens, and (2) the average number of update iterations across the entire dataset. In nearly all cases, excluding Qwen2.5-1.5B-Instruct on GSM8K with prompt 2, the token ratio does not exceed 1.1. This suggests that, unlike long CoT approaches Xu et al. (2025b), our method does not rely on generating extended outputs to achieve strong performance, thus avoiding overly verbose generation. Moreover, concerning the number of update iterations, which constitutes the core computational component of the algorithm, the average iteration counts for GSM8K and MATH-500 across various model backbones and prompts are 0.86 and 1.23, respectively. Both values are below 2, indicating that the search process in the latent space typically converges in fewer than two iterations for average-level questions. This highlights the computational efficiency and rapid convergence behavior of our method.
@@ 3.6 Qualitative Analysis
To gain deeper insight into model-generated reasoning, we analyze the distribution of the first three tokens produced under Prompt 1 in the GSM8K dataset Cobbe et al. (2021) using the Llama3.1-8B-Instruct model. The frequent appearance of nonsensical or obscure tokens—such as “thecy,” “theella,” and “Sid”—indicates that the model’s reasoning mechanisms often diverge from human cognitive strategies, even when producing superficially coherent outputs. A representative example, detailed in Table 5, further illustrates this divergence: despite generating linguistically anomalous expressions such as “total downloads of downloads,” and “down step two andLet,” the model still arrives at the correct answer. These findings suggest that the optimal reasoning trajectories for language models are fundamentally different from those of humans and may be more effectively navigated within the model’s latent space. For further discussion, see Appendix D.
Table 5: Case Studies: the questions, the answers, the CoT generations, and the LatentSeek generations are listed. The incoherence examples are marked with blue color. See Appendix D for more cases.
@@ 5 Conclusion
[…]
1Contents
Appendix A Theoretical Analysis
@@
A.1 Preliminaries: Multiple Prover Interactive Proofs and NEXP
@@
Definition A.1 (Multiple Prover Interaction).
@@
Definition A.2 ($k$-MIP Vadhan and Zhang (2002)).
@@
Definition A.3 (MIP Vadhan and Zhang (2002)).
$=\cup_{k}k-\text{MIP}$ MIP
@@
Definition A.4 (NTIME Arora and Barak (2006)).
@@
Definition A.5 (NP).
NP $=\cup_{k\in\mathbb{N}}$ NTIME($n^{k}$)
Definition A.6 (NEXP).
NEXP $=\cup_{k\in\mathbb{N}}$ NTIME($2^{n^{k}}$)
A.2 Theoretical Analysis: Independent Updating
Mentioned in Section 2, the latent representations are independent, which results in the independence of the token update process, which at first glance may constrain the model’s expressive capacity. In this subsection, we focus on decision problems Arora and Barak (2006). Despite discrepancies between theoretical frameworks and practical challenges, valuable insights can be gained. Specifically, it is possible to demonstrate that the model’s expressivity remains theoretically comparable in spite of the updating independence.
Our method is strongly related to a TCS complexity class, Multi Prover Interactive Proofs Ben-Or et al. (1988) or simply MIP. We initially outline the notations for multiple prover interactions and subsequently define the associated complexity class for our latent thought framework, which we later demonstrate is equivalent to MIP and the complexity class NEXP (solved in exponential time by a non-deterministic Turing Machine). Please refer to Section A.1 for preliminaries.
Definition A.7 (Multiple Prover Interaction).
@@ A.2 Theoretical Analysis: Independent Updating
  • Completeness: $x\in L\Rightarrow\exists P_{1},P_{2},\ldots,P_{\text{poly}(|x|)},\,Pr[\braket{(P_{1},P_{2},\ldots,P_{\text{poly}(}{x}{)}),V}{}{V}(x)=1]\geq\frac{2}{3}$
  • Soundness: $x\notin L\Rightarrow\forall P_{1},P_{2},\ldots,P_{\text{poly}(|x|)},\,Pr[\braket{(P_{1},P_{2},\ldots,P_{\text{poly}(}{x}{)}),V}{}{V}(x)=1]\leq\frac{1}{3}$
Remark A.9.
@@
Theorem A.10 establishes the efficacy of our framework. By building on the classical result of Babai et al. (1990), we derive Corollary A.11 that clearly illustrates the expressive power of our approach. Refer to Section A.3 for detailed proofs.
Corollary A.11.
NP $\subset$ NEXP $=$ MIP-Bounded
A.3 Proofs of Theorem A.10 and Corollary A.11
Theorem A.12.
@@
[…]
@@
    $Pr[\braket{(P_{1}^{\prime},P_{2}^{\prime},\ldots,P_{\text{poly}(}{x}{)}^{\prime}),V^{\prime}}{}{V^{\prime}}(x)=1]\geq\frac{2}{3}$
@@
    $Pr[\braket{(P_{1}^{\prime},P_{2}^{\prime},\ldots,P_{\text{poly}(}{x}{)}^{\prime}),V^{\prime}}{}{V^{\prime}}(x)=1]\leq\frac{1}{3}$
[…]
@@
[…]
    $Pr(\braket{(P_{1},P_{2}),V}_{V}(x)=1)$ $=Pr[\braket{(P_{1}^{\prime},P_{2}^{\prime},\ldots,P_{\text{poly}(}{x}{)}^{\prime}),V}{}{V}(x)=1]$
    $=Pr[\braket{(P_{1}^{\prime},P_{2}^{\prime},\ldots,P_{\text{poly}(}{x}{)})^{\prime},V^{\prime}}{}{V^{\prime}}(x)=1]$
@@
    $Pr[\braket{(P_{1}^{\prime},P_{2}^{\prime},\ldots,P_{\text{poly}(}{x}{)}^{\prime}),V^{\prime}}{}{V^{\prime}}(x)=1]>\frac{1}{3}$
    , this contradicts to $Pr[\braket{(P_{1}^{\prime},P_{2}^{\prime},\ldots,P_{\text{poly}(}{x}{)}^{\prime}),V^{\prime}}{}{V^{\prime}}(x)=1]\leq\frac{1}{3}$. Therefore, $\forall P_{1},P_{2}$, we have:
@@
[…]

### v2

@@ 3.2 State-of-the-art Test-time Reasoning Performance
The main experimental results are presented in Table˜1 and Table˜2.
@@
As demonstrated in Table˜2, using LLaMA3.1-8B-Instruct as the backbone, our method outperforms all baseline approaches across both the GSM8K and MATH-500 datasets. Specifically, on the GSM8K dataset, our method achieves an improvement of 14.6% points over the CoT baseline, and a 7.0%-point gain on the MATH-500 dataset. In comparison to the training-free BoN approach, our method yields improvements of 7.7% points and 3.4% points on GSM8K and MATH-500, respectively. When compared to the SFT (Magpie 25K) Xu et al. (2025a), the minimum improvement across both datasets is 12.9% points. Moreover, our approach surpasses the state-of-the-art self-reward RL-based method, Genius Xu et al. (2025a), by 5.4% points on GSM8K and 20.0 points on MATH-500. Notably, it also outperforms iCoT Deng et al. (2024)—a latent method trained on an augmented GSM8K dataset—by 36.6% points on the GSM8K test set. Even when compared to reinforcement learning with verifiable feedback methods, like SimpleRL-Zoo Zeng et al. (2025), which was trained on the base version of the model, our method exceeds it by an average of 9.1% points across both datasets. These results underscore the effectiveness of our approach in tackling reasoning tasks.
@@
AIME2024 results are listed in Table˜1. AIME2024 is a highly challenging mathematics benchmark consisting of only 30 problems, designed to evaluate advanced reasoning capabilities. Experimental results demonstrate the power and potential of our method in complex settings. As shown in Table˜1, our approach consistently outperforms all baselines, achieving an average improvement of 4.73% points over CoT across all model families and prompt configurations. Notably, even weaker models such as Mistral and Qwen2.5-1.5B-Instruct benefit from our method, with average gains of 1.67% and 3.33% points compared with CoT, respectively. The largest improvement is observed on AIME2024 (prompt 1), where our method surpasses the CoT baseline by 13.33% points and outperforms BoN by 6.67% points using Qwen2.5-7B-Instruct.
@@
As presented in Section˜3, the generalizability of our approach with respect to backbone models can be examined along two dimensions: (1) Model Families: LatentSeek demonstrates superior performance across multiple model families. Specifically, it achieves the best results when using backbones from the Qwen2, Qwen2.5, and LLaMA3.1 series. Additionally, it ranks first in four out of six evaluation settings and second in one when employing a relatively weaker backbone from the Mistral family. The suboptimal performance of the Mistral-based model is attributed to its limited capacity to function effectively as a reward model. (2) Model Size: At the 1.5B parameter scale, our method consistently outperforms all baseline models across diverse datasets and prompt types, with a particularly notable gain of 28.89% points over the CoT baseline on the GSM8K dataset (Prompt 2). At the 7–8B scale, models based on Qwen2, Qwen2.5, and LLaMA3.1 consistently surpass all baselines. In particular, Qwen2 yields an average improvement of 11.75% points over BoN on GSM8K. At the 14B scale, our approach continues to outperform all baselines, achieving a significant 6.67%-point improvement on the AIME2024 dataset. These results provide robust evidence of our method’s generalizability across diverse model families and scales.
@@ 3.3 Ideal Experiment: Perfect Sparse Reward Model
To further evaluate the effectiveness of exploration within the latent space, we conduct experiments using a Perfect Sparse Reward Model (PSRM). In this experimental setup, a reward value of 0 is assigned exclusively when the generated final answer exactly matches the ground truth. In all other cases, a reward of 1-1 is given. The results of these experiments are presented in Table 1. Despite the limited feedback provided by this sparse reward signal—particularly in contrast to the self-reward model—the PSRM demonstrates notable performance gains. Specifically, it achieves an average improvement of 19.12% score over the CoT method and surpasses the self-reward version by an average of 12.57% score. These results highlight the potential of LatentSeek, even when guided by sparse reward.
@@ 3.4 Test-Time Scaling: scaling up the iteration of LatentSeek
In addition to increasing the number of generated tokens at test time Muennighoff et al. (2025), Snell et al. (2025), we propose an alternative approach to test-time scaling: increasing the number of LatentSeek iterations. The relationship between model performance on GSM8K Cobbe et al. (2021) and the number of iterations is illustrated in Figure˜2. As shown, the ideal reward model yields a consistently improving trend and outperforms the self-reward method across all model backbones, suggesting that test-time scaling can be achieved without necessitating a dense reward function in our setting. Conversely, the self-reward method exhibits rapid initial gains followed by a plateau, attributable to the limitations in the reward model’s accuracy. Notably, with an increased number of iterations, performance eventually surpasses this plateau and continues to improve. These findings indicate that, given an appropriate reward model, searching through the latent space offers a promising new direction for test-time scaling.
@@
Following the methodology proposed by Liu et al. (2025), we conduct extreme scaling experiments using the Qwen2.5-1.5B-Instruct model Team (2024a). In contrast to Liu et al. (2025), which leverages Process Reward Models (PRM), our approach is specifically tailored to align with Outcome Reward Models (ORM). Given the current limitations in ORM design (see Section˜E.5) and the inherent challenges of self-rewarding in small-scale models, we investigate the effectiveness of this novel test-time scaling strategy through the use of the PSRM. As shown in Table˜3, Qwen2.5-1.5B-Instruct achieves a 14% point performance gain over GPT-4o on the AIME2024 benchmark. Furthermore, on the MATH-500 dataset, it attains the highest overall accuracy among all evaluated models, with a marginal 2.7% point gap behind o1-preview. Remarkably, the model requires only an average of 61.8 iterations on MATH-500 to reach the performance, outperforming a 256 compute consumption of explicit space scaling methods in terms of efficiency. This implies that, under an appropriate reward model, the latent space represents a more efficient option for test-time scaling compared to the explicit space.
@@ 3.5 Algorithmic Statistics
Table˜4 reports two key statistical features: (1) the ratio of final answer tokens to the original CoT tokens, and (2) the average number of update iterations across the entire dataset. In nearly all cases, excluding Qwen2.5-1.5B-Instruct on GSM8K with prompt 2, the token ratio does not exceed 1.1. This suggests that, unlike long CoT approaches Xu et al. (2025b), our method does not rely on generating extended outputs to achieve strong performance, thus avoiding overly verbose generation. Moreover, concerning the number of update iterations, which constitutes the core computational component of the algorithm, the average iteration counts for GSM8K and MATH-500 across various model backbones and prompts are 0.86 and 1.23, respectively. Both values are below 2, indicating that the search process in the latent space typically converges in fewer than two iterations for average-level questions. This highlights the computational efficiency and rapid convergence behavior of our method.
@@ 3.6 Qualitative Analysis
To gain deeper insight into model-generated reasoning, we analyze the distribution of the first three tokens produced under Prompt 1 in the GSM8K dataset Cobbe et al. (2021) using the Llama3.1-8B-Instruct model. The frequent appearance of nonsensical or obscure tokens—such as “thecy,” “theella,” and “Sid”—indicates that the model’s reasoning mechanisms often diverge from human cognitive strategies, even when producing superficially coherent outputs. A representative example, detailed in Table˜5, further illustrates this divergence: despite generating linguistically anomalous expressions such as “total downloads of downloads,” and “down step two andLet,” the model still arrives at the correct answer. These findings suggest that the optimal reasoning trajectories for language models are fundamentally different from those of humans and may be more effectively navigated within the model’s latent space. For further discussion, see Appendix˜H.
Table 5: Case Studies: the questions, the answers, the CoT generations, and the LatentSeek generations are listed. The incoherence examples are marked with blue color. See Appendix˜H for more cases.
@@ 5 Conclusion
Acknowledgement
We thank Yanpeng Zhao from BIGAI, Haoran Sun from Peking University for their helpful reviews of the paper.
@@ Acknowledgement
Contents
Appendix A Discussion and future works
While our work demonstrates the effectiveness of latent space optimization, we offer some discussion for inspiring future research.
Reward Models.
A primary limitation of our current approach is its reliance on a self-rewarding mechanism. While this proves effective, the optimization process is inherently constrained by the base model’s own evaluation capabilities and potential biases, lacking a truly objective external signal. The ideal scenario would involve a principled Outcome Reward Model (ORM) that assesses the final answer and provides unambiguous guidance for the search. However, as detailed in our experiments in Section˜E.5, we found that current publicly available ORMs are not yet sufficiently robust or generalizable for this task; their noisy signals proved less effective for guiding the latent space optimization than the more consistent self-reward mechanism. This highlights a critical area for future work: the development of more powerful verifiers capable of reliably scoring complex reasoning tasks.
Latent Optimization.
We adopt standard policy-gradient methods in our implementation, leaving the exploration of more advanced reinforcement learning algorithms—such as Proximal Policy Optimization (PPO)—to future work. In addition, pursuing latent-space–specific optimization may be a worth exploring direction.
Large Base Model.
Our experiments are conducted on models up to 14B-parameter scale, constrained by available computational resources. Scaling the approach to larger base models remains an important avenue for future investigation.
Appendix B Methods of Test-Time Instance-Level Reasoning
We list the formulations of two classical test-time instance-level reasoning methods:
    $\mathbf{x}^{*}=\operatorname*{arg\,max}_{\mathbf{x}}\pi(\mathbf{x}\mid\mathbf{c})$ (9)
  • Best-of-N (BoN): Given $N$ i.i.d sequences $\mathbf{x}_{(1)},\mathbf{x}_{(2)},\ldots,\mathbf{x}_{(N)}\sim\pi(\cdot\mid\mathbf{c})$,
    $\mathbf{x}^{*}=\operatorname*{arg\,max}_{\mathbf{x}\in\{\mathbf{x}_{(1)},\mathbf{x}_{(2)},\ldots,\mathbf{x}_{(N)}\}}R(\mathbf{x},c)$ (10)
Appendix C Theoretical Analysis
@@
C.1 Preliminaries: Multiple Prover Interactive Proofs and NEXP
@@
Definition C.1 (Multiple Prover Interaction).
@@
Definition C.2 ($k$-MIP Vadhan and Zhang (2002)).
@@
Definition C.3 (MIP Vadhan and Zhang (2002)).
$\text{MIP}=\cup_{k}k-\text{MIP}$
@@
Definition C.4 (NTIME Arora and Barak (2006)).
@@
Definition C.5 (NP).
NP $\cup_{k\in\mathbb{N}}$ NTIME($n^{k}$)
Definition C.6 (NEXP).
NEXP $\cup_{k\in\mathbb{N}}$ NTIME($2^{n^{k}}$)
C.2 Theoretical Analysis: Independent Updating
Mentioned in Section˜2, the latent representations are independent, which results in the independence of the token update process, which at first glance may constrain the model’s expressive capacity. In this subsection, we focus on decision problems Arora and Barak (2006). Despite discrepancies between theoretical frameworks and practical challenges, valuable insights can be gained. Specifically, it is possible to demonstrate that the model’s expressivity remains theoretically comparable in spite of the updating independence.
Our method is strongly related to a TCS complexity class, Multi Prover Interactive Proofs Ben-Or et al. (1988) or simply MIP. We initially outline the notations for multiple prover interactions and subsequently define the associated complexity class for our latent thought framework, which we later demonstrate is equivalent to MIP and the complexity class NEXP (solved in exponential time by a non-deterministic Turing Machine). Please refer to Section˜C.1 for preliminaries. We list the relationship as follows:
  • MIP: A verifier interacts with multiple non-communicating provers to decide membership in a language.
  • LatentSeek: the optimized tokens, as they are updated independently, act like the provers who jointly determine the outcome.
  • MIP-Bounded: Different from MIP, each prover in LatentSeek can only generate 1 token, whose number of bits is bounded. Furthermore, a polynomial proportion of tokens is allowed to be optimized, leading to a polynomial number of provers.
Definition C.7 (Multiple Prover Interaction).
@@ C.2 Theoretical Analysis: Independent Updating
  • Completeness: $x\in L\Rightarrow\exists P_{1},P_{2},\ldots,P_{\text{poly}(|x|)},\,Pr[\braket{(P_{1},P_{2},\ldots,P_{\text{poly}(}{x}{)}),V||V}(x)=1]\geq\frac{2}{3}$
  • Soundness: $x\notin L\Rightarrow\forall P_{1},P_{2},\ldots,P_{\text{poly}(|x|)},\,Pr[\braket{(P_{1},P_{2},\ldots,P_{\text{poly}(}{x}{)}),V||V}(x)=1]\leq\frac{1}{3}$
Remark C.9.
@@
Theorem˜C.10 establishes the efficacy of our framework. By building on the classical result of Babai et al. (1990), we derive Corollary˜C.11 that clearly illustrates the expressive power of our approach. Refer to Section˜C.3 for detailed proofs.
Corollary C.11.
NP $\subset$ NEXP == MIP-Bounded
C.3 Proofs of Theorem˜C.10 and Corollary˜C.11
Theorem C.12.
@@
We will first prove that $\text{MIP}\subseteq\text{MIP-Bounded}$.
@@
    $Pr[\braket{(P_{1}^{\prime},P_{2}^{\prime},\ldots,P_{\text{poly}(}{x}{)}^{\prime}),V^{\prime}||V^{\prime}}(x)=1]\geq\frac{2}{3}$
@@
    $Pr[\braket{(P_{1}^{\prime},P_{2}^{\prime},\ldots,P_{\text{poly}(}{x}{)}^{\prime}),V^{\prime}||V^{\prime}}(x)=1]\leq\frac{1}{3}$
Therefore, $L\in\text{MIP-Bounded}$, and thus $\text{MIP}\subseteq\text{MIP-Bounded}$.
Next, we are going to prove $\text{MIP-Bounded}\subseteq\text{MIP}$.
$\forall L\in\text{MIP-Bounded}$, $\forall x$, in the $j$-th turn, the proofs offer by the bounded provers are denoted as: $a^{\prime}_{1j},a^{\prime}_{2j},a^{\prime}_{3j},\ldots,a^{\prime}_{\text{poly}(|x|),j}$, the verifier as $V^{\prime}$. We design $V$ as follows: for each concat string $s=(x,q_{11},a_{11},\ldots,)$, $V$ first truncates the first prover’s answer to length of $\text{LIM}_{1}:=C\cdot\lfloor\frac{\text{poly}(|x|)}{2}\rfloor$ bits and the second prover’s answer to length of $\text{LIM}_{2}:=C\cdot(\text{poly}(|x|)-\lceil\frac{\text{poly}(|x|)}{2}\rceil+1)$ bits. The resulted string is denoted as $s^{\prime}:=(x,q_{11},a_{11;0:\text{LIM}_{1}},q_{21},a_{21;0:\text{LIM}_{2}}\ldots)$, and output $V^{\prime}(s^{\prime})$.
@@
    As $V$ first truncate answer, therefore, for $j$-turns, we can truncate $a_{1j}$ to $a_{1j;0:\text{LIM}_{1}}$ and $a_{2j}$ to $a_{2j;0:\text{LIM}_{2}}$, and follow the division process depicted in the first case of proving $\text{MIP}\subseteq\text{MIP-Bounded}$ to gain $P_{1}^{\prime},P_{2}^{\prime},P_{3}^{\prime},\ldots,P_{\text{poly}(|x|)}^{\prime}$ provers. Thus:
    $Pr(\braket{(P_{1},P_{2}),V}_{V}(x)=1)$ $=Pr[\braket{(P_{1}^{\prime},P_{2}^{\prime},\ldots,P_{\text{poly}(}{x}{)}^{\prime}),V||V}(x)=1]$
    $=Pr[\braket{(P_{1}^{\prime},P_{2}^{\prime},\ldots,P_{\text{poly}(}{x}{)})^{\prime},V^{\prime}||V^{\prime}}(x)=1]$
@@
    $Pr[\braket{(P_{1}^{\prime},P_{2}^{\prime},\ldots,P_{\text{poly}(}{x}{)}^{\prime}),V^{\prime}||V^{\prime}}(x)=1]>\frac{1}{3}$
    , this contradicts to $Pr[\braket{(P_{1}^{\prime},P_{2}^{\prime},\ldots,P_{\text{poly}(}{x}{)}^{\prime}),V^{\prime}||V^{\prime}}(x)=1]\leq\frac{1}{3}$. Therefore, $\forall P_{1},P_{2}$, we have:
@@
Hence, we have $L\in\text{MIP}$, which means that $\text{MIP-Bounded}\subseteq\text{MIP}$.